# SplineSketch: Even More Accurate Quantiles with Error Guarantees

### v1

@@ 3.1. Consolidating buckets
In fact, the consolidate method allows to easily change the number $k$ of buckets, thus making the sketch larger or smaller. Changing $k$ to $k^{\prime}$ buckets is done in a straightforward way: If we increase the sketch size, i.e., $k^{\prime}>k$, we perform $k^{\prime}-k$ splits without any join, first dealing with buckets with exceeded bound and then in the order of the heuristic error. On the other hand, if $k^{\prime}<k$, after performing the necessary splits due to exceeding the bound, we execute $k-k^{\prime}$ joins in the order of heuristic error, on bucket pairs satisfying Definition 1 and without performing any splits. These separate splits or joins may be followed by performing consolidation in a normal way, where we always join after splitting and vice versa. If $k^{\prime}$ is substantially different from $k$, say $|k^{\prime}-k|>0.25\cdot k$, we also end the current epoch, i.e., reset the protection bit vector. Note that our analysis in Appendix B does not account for sketch resizing.
@@ 3.2. Merging Two SplineSketches
We merge two sketches $S_{1}$ and $S_{2}$ with $k_{1}$ and $k_{2}$ buckets, respectively, as follows. Without loss of generality, at least as many items have been inserted into $S_{1}$ as into $S_{2}$. First, we create a buffer that contains all the items from the buffers of the two sketches we are merging.
@@
Second, we take the union of all the thresholds $\tau$ from both sketches. After removing duplicates, this gives up to $k_{1}+k_{2}$ thresholds. For each of these thresholds $\tau$, we query for its rank in both sketches and add these two ranks to get the rank of $\tau$ in the merged sketch. We set the counter of each bucket to be the difference between the ranks of the endpoints.
Third, the new sketch inherits the bit vector for threshold protection from $S_{1}$ as it summarizes a larger input than $S_{2}$. If the total input size of the new sketch is larger than the epoch end for $S_{1}$, we set all thresholds to the unprotected state, starting a new epoch.
@@
Finally, we run the consolidate method that merges a potentially overflown buffer to buckets and more importantly, reduces the number of buckets to $k_{1}$. Namely, we repeatedly take a joinable pair of adjacent buckets whose joining would result in a bucket with the lowest heuristic error and join it without any split before we have $k_{1}$ buckets.
3.3. Dealing with High-Frequency Items
The algorithm as described above works and satisfies the worst-case bound as long as all items’ frequencies are below $n/k$.
There are two ways to deal with high-frequency items. First, one can satisfy the assumption of no high-frequency items on any input by running the Misra-Gries (MG) sketch (Misra and Gries, 1982) of size $O(k)$ in parallel. That is, we first insert an incoming item $x$ into the MG sketch. If $x$ is stored by MG, then we do not add $x$ to SplineSketch; otherwise, we add $x$ to SplineSketch. The latter case may result in some items being ousted from the MG sketch, which we add to SplineSketch together with their frequencies in MG.
However, using MG or another heavy-hitter algorithm would make our sketch more complicated and slower. In the actual implementation, we do not maintain a heavy-hitter sketch in parallel, but rather try to recognize heuristically whether or not a bucket represents a single item of high frequency as it has a very small relative length. See Section 4.2 for details.
@@ 4. Implementation Details
So far, we focused on answering rank queries. In order to implement a quantile query, i.e., returning an estimated $r$-th smallest number for a given $r$, we invert the interpolation, and query the inverse function at $r$; the inverse exists since buckets are non-empty and the PCHIP interpolation is thus strictly increasing from $\tau_{1}$ to $\tau_{k}$. The inverse function can be numerically evaluated by using, for example, the Newton’s method or even just binary search. Note that our algorithm, similarly as q-digest, may return a number that did not appear in the stream, unlike comparison-based methods such as the KLL sketch (Karnin et al., 2016).
4.1. Implementations of the consolidate method
We describe two particular implementations of the consolidate method. The first achieves a low amortized update time, while the second is easier to implement and works well in practice.
@@ 4.1. Implementations of the consolidate method
To obtain asymptotically fast consolidate method, one can use a combination of doubly-linked lists for maintaining buckets and two heaps, sorted by the bucket size and heuristic error, respectively. This approach is formalized in the following lemma. Since the buffer size is $O(k)$, we get that the amortized complexity of an update with heap-based consolidate is $O(\log k)$.
@@
lemmaheapBasedConsolidate Consolidation can be implemented in $O(k\log k)$ time.
@@
We maintain buckets in a doubly linked list $L$. We use a queue $Q_{S}$ to keep track of buckets that need to be split due to violating bucket bound. We also use two heaps, $H_{S}$ and $H_{J}$, to keep track of buckets that might get split due to heuristic error and joinable pairs of buckets, respectively. $H_{S}$ is a max-heap with heuristic errors as keys, and $H_{J}$ is a min-heap with the keys being the heuristic errors after the join. In both heaps, every node contains a pointer to the corresponding bucket (or pair of buckets) in $L$. Conversely, every bucket in $L$ contains a (constant size) list of pointers to nodes of $H_{S}$ and $H_{J}$ containing that bucket. Both heaps can be initialized at the beginning of consolidate in $O(k)$ time.
After every split and join performed by consolidate method, we update the $Q_{S},H_{S}$, and $H_{J}$ according to the new bucket sizes. Since every split and join affect only the sizes of constant number of adjacent buckets, the update can be done in $O(1)$ time using the pointers that we maintain. The modifications of the keys and deletions of arbitrary nodes in the heaps can be done in $O(\log k)$ time using standard decrease/increase key operation.
Every split introduces a constant number of new protected thresholds and every join is accompanied with a corresponding split. Since the number of possible protected thresholds is bounded by $O(k)$, the total number of splits and joins performed by consolidate is also $O(k)$. Therefore, the total time complexity of the heap based implementation of consolidate is $O(k\log k)$.
@@
In our implementation, we use practically more efficient consolidate method without the need for linked lists and heaps. We process the buckets in iterations such that each iteration splits or joins one bucket at most once and not both at the same time; note that more iterations are possibly needed as one bucket may be split or joined mutlple times during consolidation. We stop when the last iteration does not change any threshold. Throughout this process, we keep the buffer and the old buckets for more accurate setting of counters for the new buckets.
@@
In each iteration, we first collect the buckets with size exceeding the bound (1) into a set $E$. We also keep a list of splittable buckets (Def. 2), sorted non-increasingly by their heuristic error, and a list of joinable pairs of buckets (Def. 1), sorted non-decreasingly by their heuristic errors after potentially performing their join. We select $|E|$ joinable pairs according to their order that do not overlap buckets in $E$ or a previously selected joinable pairs. Next, from the remaining buckets (not yet split or joined), we consider the splittable bucket with the largest heuristic error $h_{\mathrm{split}}$ (if any) and the joinable pair with the lowest heuristic error $h_{\mathrm{join}}$ and if $h_{\mathrm{split}}>1.5\cdot h_{\mathrm{join}}$, then we perform the split and join of these buckets. After collecting non-overlapping sets of buckets to split and join, we perform these splits and joins in one pass over the buckets, using the buffer and interpolation over the buckets before the consolidation; the usage of buckets before consolidation is important otherwise error may accumulate.
@@
Finally, we remark that the multiplicative constant $C_{b}$ in the bucket bound (1) resulting from the analysis in App. B is too large for practical usage, and we just use $C_{b}=3$, increasing $C_{b}$ during an epoch if needed (i.e., if there are no joinable pairs when a bucket must be split due to exceeding the bound) and resetting it to $C_{b}=3$ at the end of each epoch; we note that such increases of $C_{b}$ happen rarely in practice.
4.2. Handling Frequent Items without a Heavy-Hitter Sketch
@@
The potential presence of high-frequency items makes the implementation more convoluted. We chose not to use a separate heavy-hitter sketch to deal with frequent items as it would make SplineSketch slower when processing the stream while item frequencies are relatively small in many applications. Instead, we set a lower limit on the relative length of a bucket, depending on numerical precision $\delta$ of data. That is, a bucket $(\tau_{i},\tau_{i+1}]$ must have length $\tau_{i+1}-\tau_{i}\geq\delta^{\prime}\cdot\max\{|\tau_{i}|,|\tau_{i+1}|,\varepsilon\}$, where $\delta^{\prime}>\delta$ is a parameter and $\varepsilon>0$ is the smallest non-zero absolute value of an input item ($\varepsilon$ is intended to avoid too small buckets around 0). In our prototype implementations in Python and in Java with the double data type, we set $\delta^{\prime}=10^{-8}$. We do not split a bucket that would violate this relative length bound, even if it does not satisfy the size bound (1); this way, we avoid too many splits of a bucket with a frequent item.
@@
We also deal with repeated thresholds when creating initial buckets by equally spaced selection (after the buffer gets full for the first time). Namely, we remove duplicate thresholds and instead, add new thresholds relatively close to the previously repeated thresholds (but not too close to violate the relative length bound) so that we end up with $k$ thresholds in total. This makes our approach robust to frequent items that appear from the beginning: The new thresholds will in fact heuristically guarantee that the frequent item gets its own bucket if it is frequent from the very beginning.
@@ 5. Experimental Evaluation
We have implemented SplineSketch as a prototype in Python and Java. Here, we provide a comparison with state-of-the-art quantile sketches for the uniform rank error with an available implementation, namely, $t$-digest (Dunning and Ertl, 2019), MomentSketch (Gan et al., 2018), and KLL (Karnin et al., 2016); see Appendix C for details. We run the sketches on three real-world and ten synthetically generated datasets (App. C). We measure the average and maximum absolute rank errors for $10^{5}$ evenly spaced queries, together with average update and query times. The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers.
(a) Uniform distribution
(b) Pareto distribution
(c) Gumbel distribution
(d) Loguniform distribution
(e) Normal distribution with a large parameter change after $n/2$ items
(f) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 3. Average rank error (log-scale) on synthetic datasets with $n=10^{8}$ items depending on the sketch size in bytes. The average is taken over $10^{5}$ evenly spaced queries.

### v2

@@ 12   Mark thresholds $\tau_{i-1},\tau^{\prime},\tau_{i}$ as protected;
Bucket consolidation allows to change the number $k$ of buckets, making the sketch larger or smaller as desired. Changing from $k$ to $k^{\prime}$ buckets is straightforward: If we increase the sketch size, i.e., $k^{\prime}>k$, we perform $k^{\prime}-k$ splits without any join, first dealing with buckets exceeding bound (1) and then in the order of the heuristic error. On the other hand, if $k^{\prime}<k$, after performing the necessary splits due to exceeding the bound (and the corresponding joins), we execute $k-k^{\prime}$ joins in the order of their heuristic errors after joining, on bucket pairs satisfying Definition 1 and without performing any splits. These separate splits or joins may be followed by performing Consolidate in a normal way, where we always join after splitting. If $k^{\prime}$ is substantially different from $k$, say $|k^{\prime}-k|>0.25\cdot k$, we also reset the protection bit vector.
@@ 3.2. Merging Two SplineSketches
We describe how to merge two SplineSketches $S_{1}$ and $S_{2}$ with $k_{1}$ and $k_{2}$ buckets, respectively, into one sketch $S$ for the combined datasets. We merge one sketch into the other, namely $S_{2}$ into $S_{1}$, but the user can specify whether to merge $S_{2}$ into $S_{1}$, or vice versa. The default choice is that we merge the sketch summarizing a smaller number of items into the one with more items. Furthermore, in the analysis, we assume that $k_{1}=k_{2}$ but the merge operation works even for $k_{1}\neq k_{2}$.
@@
First, we create a buffer that contains all the items from the buffers of the two sketches we are merging. Second, we take the union of all the thresholds $\tau$ from both sketches. After removing duplicates, this gives up to $k_{1}+k_{2}$ thresholds. For each of these thresholds $\tau$, we query for its rank in both input sketches and add these two ranks to get the rank of $\tau$ in the merged sketch $S$. We set the counter of each bucket to be the difference between the ranks of the endpoints. Third, the new sketch inherits the bit vector for threshold protection from $S_{1}$, meaning that any threshold taken from $S_{1}$ remains protected if it was protected in $S$. However, if the total input size of the new sketch is larger than the epoch end for $S_{1}$, we set all thresholds to the unprotected state, starting a new epoch.
@@
Finally, we run Consolidate that reduces the number of buckets to $k_{1}$ and also merges the buffer to buckets. Namely, we repeatedly take a joinable pair of adjacent buckets whose joining would result in a bucket with the lowest heuristic error and join it without any split until we have $k_{1}$ buckets.
3.3. Theoretical Properties
We now provide error bounds that hold for any input in the mergeability setting, which captures the streaming model as we can interpret stream updates as merges with a single-item sketch.
Theorem 4 (See Appendix B.3).
Suppose that we build a SplineSketch using any sequence of pairwise merge operations executed on $n$ data items, with every intermediate SplineSketch consisting of $k$ buckets. Then the resulting SplineSketch has a worst-case rank error of $O(\log(\alpha)\cdot n/k)$.
The analysis assumes that no element appears more than $n/k$ times. In order to lift this assumption, we can run a heavy-hitter Misra-Gries (MG) sketch (Misra and Gries, 1982) of size $O(k)$ in parallel. We describe how to effectively combine it with SplineSketch in Sec. 4.2.
We note that state-of-the-art quantile sketches achieve better worst-case accuracy-space trade-offs: Namely, KLL using space of $O(k)$ memory words has worst-case rank error $O(n/k)$ with constant probability (Karnin et al., 2016). The deterministic Greenwald-Khanna (GK) sketch guarantees rank error $O(n/k)$ but with worst-case space $O(k\cdot\log(n/k))$ (Greenwald and Khanna, 2001); the $\log(n/k)$ factor is in general incomparable to the $\log(\alpha)$ in the error for SplineSketch. However, both KLL and GK achieve significantly worse error in practice, as we show in Sec. 5. As for mergeability, KLL is fully mergeable while retaining the guarantees (Karnin et al., 2016) but GK is known to be only one-way mergeable (Agarwal et al., 2013). Other quantile sketches have worse rank error or do not provide a bounded rank error; see Sec. 2.
Our analysis is very flexible with respect to individual components and parameters of the sketch, with only constant factor changes (such as the value of $C_{b}$ in Eqn. 1). Namely, it is independent of how the heuristic error is computed (and of the value of $\gamma$ in Def. 2) and as for other parameters, the analysis needs that epoch lengths increase geometrically. As for the interpolation, it can be changed to linear or other monotone interpolation that is exact on the thresholds if it satisfies that in the middle of a non-empty bucket, the interpolation is bounded away from the minimum and maximum estimated values inside the bucket, i.e., a bucket counter is divided somewhat evenly during a split (a property similar to Lemma 1 in Appendix A).
Regarding the effect of resizing the sketch from $k^{\prime}$ buckets to $k^{\prime\prime}$ buckets on the guarantees, the error bound of $O(\log(\alpha)\cdot n/k)$ holds with $k=\min\{k^{\prime},k^{\prime\prime}\}$ after performing the resizing. To the best of our knowledge, resizability is not addressed in the literature on quantile summaries yet.
As for the time complexities, the time for one rank or quantile query is $O(\log k)$, as derived at the beginning of Sec. 3. The amortized update time can be made $O(\log k)$ by a heap-based implementation of Consolidate, as explained in Sec. 4.1. Finally, the merge time is $O(k\cdot\log k)$, again for the heap-based Consolidate. These theoretical time complexities are asymptotically the same as for KLL, GK, and $t$-digest, albeit particular implementations of the sketches may not satisfy the theoretical bounds.
@@ 4. Implementation Details
Here, we explain the details of our SplineSketch implementation (in Python and Java) and elaborate on the update time.
First, we naturally use the double data type (64-bit floating point numbers) for thresholds. This, however, implies some limitations, such as worse accuracy for some inputs with many significant digits (e.g., the SOSD benchmark (Kipf et al., 2019; Marcus et al., 2020)), which is due to the fact that many different items get rounded to the same double value. This can be addressed by changing the underlying data type for thresholds, e.g., to 128-bit floating-point numbers or to unsigned 64-bit integers. Note that the internal PCHIP interpolation would still need to use floating-point numbers.
@@
So far, we focused on answering rank queries. In order to implement a quantile query, i.e., returning an estimated $r$-th smallest number for a given $r$, we invert the interpolation, and query the inverse function at $r$; the inverse exists since buckets are non-empty and the PCHIP interpolation is thus strictly increasing from $\tau_{1}$ to $\tau_{k}$. The inverse function can be numerically evaluated by using, for example, the Newton’s method or even just binary search. Note that our algorithm, similarly as $q$-digest (Shrivastava et al., 2004), may return a number that did not appear in the stream, unlike comparison-based methods such as the KLL sketch (Karnin et al., 2016) or the Greenwald-Khanna sketch (Greenwald and Khanna, 2001).
4.1. Implementations of Bucket Consolidation
We describe two particular implementations of Consolidate. The first achieves a low amortized update time, while the second is easier to implement and works well in practice.
@@ 4.1. Implementations of Bucket Consolidation
To obtain an asymptotically fast Consolidate, one can use a combination of doubly-linked lists for maintaining buckets and two heaps, sorted by the bucket size and heuristic error, respectively. This approach is formalized in the following lemma. Since the buffer size is $O(k)$, we get $O(\log k)$ amortized complexity of an update with heap-based Consolidate.
@@
lemmaheapBasedConsolidate Consolidate can be implemented in $O(k\log k)$ time.
@@
We maintain buckets in a doubly linked list $L$. We use a queue $Q_{S}$ to keep track of buckets that need to be split due to violating bucket bound. We also use two heaps: a max-heap $H_{S}$ to keep track of buckets that might get split due to the heuristic error (2), ordered by their heuristic errors, and a min-heap $H_{J}$ for joinable pairs of buckets, ordered by their heuristic errors after temporarily joining each pair. In both heaps, every node contains a pointer to the corresponding bucket (or pair of buckets) in $L$. Conversely, every bucket in $L$ contains a (constant size) list of pointers to nodes of $H_{S}$ and $H_{J}$ containing that bucket. Both heaps can be initialized in $O(k)$ time.
After every split and join performed by Consolidate, we update $Q_{S},H_{S}$, and $H_{J}$ according to the new bucket sizes. Since every split and join affects only the sizes and heuristic errors of a constant number of adjacent buckets, the update can be done in $O(1)$ time using the pointers that we maintain. The modifications of the keys and deletions of arbitrary nodes in the heaps can be done in $O(\log k)$ time using standard decrease/increase key operation.
Every split introduces a constant number of new protected thresholds and every join is accompanied with a corresponding split. Since the number of possible protected thresholds is bounded by $O(k)$, the total number of splits and joins performed by Consolidate is also $O(k)$. Therefore, the total time complexity of the heap based implementation of Consolidate is $O(k\log k)$. ∎
@@
In our implementation, we use a practically more efficient Consolidate without the need for linked lists and heaps. We process the buckets in iterations such that each iteration splits or joins one bucket at most once and not both at the same time; note that more iterations are possibly needed as one bucket may be split or joined multiple times during Consolidate. We stop when the last iteration does not change any threshold. Throughout this process, we keep the buffer and the old buckets for more accurate setting of counters for the new buckets.
@@
In each iteration, we first collect the buckets with size exceeding the bound (1) into a set $E$. We also keep a list of splittable buckets (Def. 2), sorted non-increasingly by their heuristic error, and a list of joinable pairs of buckets (Def. 1), sorted non-decreasingly by their heuristic errors after temporarily performing their joins. We select $|E|$ joinable pairs according to their order that do not overlap buckets in $E$ or a previously selected joinable pairs. Next, from the remaining buckets (not yet split or joined), we consider the splittable bucket with the largest heuristic error $h_{\mathrm{split}}$ (if any) and the joinable pair with the lowest heuristic error $h_{\mathrm{join}}$ and if $h_{\mathrm{split}}>\gamma\cdot h_{\mathrm{join}}$ for $\gamma>1$, then we perform the split and join of these buckets; in our implementation, we use $\gamma=1.5$. After collecting non-overlapping sets of buckets to split and join, we perform these splits and joins in one pass over the buckets, using the buffer and interpolation over the buckets in their state before executing Consolidate to estimate ranks in the middle of split buckets; the usage of former buckets is important to decrease the error accumulation. We refer to the prototypes referenced in Sec. 5 for details.
@@
Finally, we remark that the multiplicative constant $C_{b}$ in the bucket bound (1) resulting from the analysis in App. B is too large for practical usage, and we just use $C_{b}=3$, while increasing $C_{b}$ during an epoch if needed and resetting it to $C_{b}=3$ at each epoch end. That is, if there are no joinable pairs when a bucket must be split due to exceeding the bound, we double $C_{b}$ and reset the iteration; we note that such increases of $C_{b}$ happen rarely in practice.
4.2. Handling Frequent Items
The theoretical error guarantees of SplineSketch, outlined in Sec. 3.3, require a method for filtering items with frequency above $n/k$. This can be done using a heavy hitter sketch such as the Misra-Gries (MG) summary (Misra and Gries, 1982). It is also important in practice, as our experiments in Sec. 5.3 demonstrate. However, using the MG sketch comes at a cost of increased update, merge, and query times, and therefore, we provide two versions of SplineSketch, with and without MG.
SplineSketch with MG
The MG sketch stores up to k1k-1 items, each with a counter $c_{x}$. When a new element $x$ arrives, if $x$ is already in the sketch, its counter is incremented. Otherwise, if $x$ is not in the sketch, we add it with a counter of $1$. If adding a new item causes the sketch to reach its capacity of $k$ items, all counters are decremented. Any item with a counter that hits zero is then removed. It can be shown that the counters in the sketch underestimate the true frequencies by at most $n/k$. In particular, any element with a frequency greater than $n/k$ is present in the sketch (Misra and Gries, 1982). The MG sketch is also fully mergeable (Agarwal et al., 2013).
In our implementation, each item $x$ in the MG sketch has an additional counter, $\hat{c}_{x}$, which is never decreased. It represents the number of times $x$ has been inserted since the beginning of the stream or its last removal from the MG sketch. During Consolidate, the MG sketch processes a buffer of new items in two stages. First, we filter out buffer items that are present in MG and increase their sketch counters by the buffer frequencies. For the remaining distinct items in the buffer, we process them one by one, and add them to the sketch with both their counters ($c_{x}$ and $\hat{c}_{x}$) set to their buffer frequency. If at any point the MG sketch becomes full, we decrease the $c_{x}$ counters of all items in the MG sketch by the smallest counter. We then remove all items with $c_{x}=0$ and add them back into the buffer with frequency $\hat{c}_{x}$. Finally, the items that remain in the buffer are processed by standard Consolidate.
Last, after processing the whole input, the extra space used by the MG sketch is compressed by removing all the items $x$ with frequency $\hat{c}_{x}<n/(2\cdot k)$. Additionally, if $\ell$ is the number of remaining items with $\hat{c}_{x}\geq n/(2\cdot k)$, we resize SplineSketch so that there are $\max\{k-\ell,k/2,6\}$ buckets. Thus, if $\ell\leq k/2$, there will be $k$ buckets and MG items in total. Based on the standard analysis of the Misra-Gries algorithm (Misra and Gries, 1982), each input element is inserted into the SplineSketch at most $n/k$ times. To process a rank query $y$ in the MG sketch, we sum the frequencies $\hat{c}_{x}$ for all items $x\leq y$. This sum is not an approximation with respect to the items summarized by the MG sketch, so it does not increase the error. We then add the result of the SplineSketch query for $y$, i.e., the $y$’s rank estimate based on the buckets and the buffer.
SplineSketch without MG
Since item frequencies can be relatively small in many applications, we also implement a faster and simpler version without heavy-hitter filtering. Still, this version tries to apply a heuristic detection of buckets with high-frequency items.
@@ 4.2. Handling Frequent Items
Namely, we set a lower limit on the relative length of a bucket, depending on numerical precision $\delta$ of data. That is, a bucket $(\tau_{i},\tau_{i+1}]$ must have length $\tau_{i+1}-\tau_{i}\geq\delta^{\prime}\cdot\max\{|\tau_{i}|,|\tau_{i+1}|,\zeta\}$, where $\delta^{\prime}>\delta$ is a parameter and $\zeta>0$ is the smallest non-zero absolute value of an input item ($\zeta$ is intended to avoid too small buckets around 0). In our prototype implementations with the double data type, we set $\delta^{\prime}=10^{-8}$. We do not split a bucket that would violate this relative length bound, even if it does not satisfy the size bound (1); this way, we avoid too many splits of a bucket with a frequent item.
@@
We also deal with repeated thresholds when creating initial buckets in Alg. 1, by removing thresholds that would violate the bucket length bound, replacing them by new thresholds relatively close to the previous threshold so that the resulting buckets would have length close to the relative length bound and we end up with $k$ thresholds. This in fact heuristically guarantees that a frequent item gets its own bucket if it is frequent from the beginning.
@@ 5. Experimental Evaluation
We have implemented SplineSketch as a prototype in Python and Java and compared the Java version with state-of-the-art quantile sketches for the uniform rank error with an available implementation, namely, $t$-digest (Dunning and Ertl, 2019), MomentSketch (Gan et al., 2018), and KLL (Karnin et al., 2016). Our implementation of SplineSketch and the code to run the experiments or generate the synthetic datasets, is available at supplementary repository (https://github.com/PavelVesely/SplineSketch-experiments). We measure the average and maximum absolute rank errors for $10^{5}$ evenly spaced queries, together with average update (or merge if applicable) and query times. Experiments were performed on an AMD EPYC 7302 (3 GHz) server with 251 GB RAM and SSD disk, and run in memory, with the datasets, queries, and sketches’ output stored on disk.
Quantile sketches.
We have used the Java prototype of SplineSketch, in two versions: with and without the Misra-Gries sketch. We have also evaluated $t$-digest (Dunning and Ertl, 2019) (v3.3, https://github.com/tdunning/t-digest), MomentSketch (Gan et al., 2018) (https://github.com/stanford-futuredata/msketch), and KLL (Karnin et al., 2016) (Java implementation by Apache DataSketches, v6.0.0, https://github.com/apache/datasketches-java), and GKAdaptive, a practically well-performing variant of the Greenwald-Khanna sketch (Greenwald and Khanna, 2001), implemented in Java according to (Luo et al., 2016). For $t$-digest, we use the default merging variant with $k_{0}$ scale function that aims at the uniform error (Dunning and Ertl, 2019). KLL was used in the variant with the double data type, i.e., an instance of KllDoublesSketch. We note that the MomentSketch implementation only answers quantile queries, while our experiments required rank estimates; we have simulated rank queries using binary search that was done for all $m$ queries in parallel in order to decrease the number of calls to the quantile query method from $\approx m\log_{2}n$ (querying just one rank) to $\approx\log_{2}n$ calls (querying $m$ ranks), to avoid repeated computation of the distribution fitting the moments and log-moments, which is a costly operation. We have not included other quantile sketches because they have worse error than KLL or t-digest, lack available implementations (e.g., algorithms from (Schiefer et al., 2023; Gribelyuk et al., 2025)), or aim for different error guarantees; namely, ReqSketch (Cormode et al., 2023) provides relative rank error and its space usage is much worse than KLL’s, while the algorithmic technique is similar. DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) aim for the relative value error, which captures the distribution tails but unlike the rank error, fails to provide a reasonable approximation for most of the distribution; indeed, we have verified that the uniform rank error of DDSketch is 10-100 times worse than KLL and even more compared to SplineSketch; see the supplementary repository for details.
The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers. Namely, for SplineSketch and $t$-digest we count 16 bytes per bucket/centroid, for MomentSketch we measured $16k+16$ bytes ($k$ moments and log-moments, minimum, and maximum), for KLL we count 8 bytes per stored item, and for GKAdaptive, we count 24 bytes per a stored tuple of one item and its rank lower and upper bounds. For SplineSketch with MG, we also account for the MG sketch size, after removing non-frequent items from the MG sketch and reducing the number of buckets based on the number of heavy hitters as described in Sec. 4.2.
Datasets.
We use three real-world datasets, the first two from UCI Machine Learning Repository (Kelly et al., 2023) : the HEPMASS dataset (Whiteson, 2016) (n=10 500 000n=10\,500\,000 items from the 2nd column), the Power dataset (Hebrail and Berard, 2006) (n=2 075 259n=2\,075\,259 items from the 3rd column), and the Books dataset from SOSD (Kipf et al., 2019; Marcus et al., 2020). We have synthetically generated 7 datasets by drawing i.i.d. (independent and identically distributed) samples from a range of distributions: uniform, normal, Pareto, Gumbel, lognormal, loguniform, and randomly signed loguniform. We have also generated three datasets with distribution change, all starting with $n/2$ items from a normal distribution and then either changing the parameters of the normal distribution (two options for the parameter change) or adding samples from a set of 42 distinct items, which will then all have high frequency. Finally, we generated a sorted input with frequent items, each with with random frequency between 1 and $n/50$. All synthetic datasets for accuracy experiments consist of $n=10^{8}$ items.
Rank queries are generated by equally spaced selection from sorted data and are executed in one batch whenever possible. While all of the sketches allow any number to be queried, using only the data points for queries leads to the same results.
5.1. Evaluation in the Streaming Setting